Consider the following 2017 data for several pharmaceutical firms ($ millions):

Average Current
Liabilities
Cash from
Operations
Expenditures
on PPE
Merck & Co., Inc. $17,909 $6,447 $1,888
Pfizer Inc. 30,771 16,470 1,956
Abbott Laboratories 7,786 5,570 1,135
Johnson & Johnson 28,412 21,056 3,279


a. Compute the operating cash flow to current liabilities (OCFCL) ratio for each firm.
(Round your answer to two decimal places.)

Merck & Co., Inc.
Pfizer Inc.
Abbott Laboratories
Johnson & Johnson


b. Compute the free cash flow for each firm.

Merck & Co., Inc.
Pfizer Inc.
Abbott Laboratories
Johnson & Johnson
Answer #1

a. Compute the operating cash flow to current liabilities (OCFCL) ratio for each firm.
(Round your answer to two decimal places.)

Merck & Co., Inc. 6447/17909 = 0.36
Pfizer Inc. 16470/30771 = 0.54
Abbott Laboratories 5570/7786 = 0.72
Johnson & Johnson 21056/28412 = 0.74


b. Compute the free cash flow for each firm.

Merck & Co., Inc. 6447-1888 = 4559
Pfizer Inc. 16470-1956 = 14514
Abbott Laboratories 5570-1135 = 4435
Johnson & Johnson 21056-3279 = 17777